OpenStudy (anonymous):

Solve the system of equations. x-4y=5 x=3y-2

OpenStudy (anonymous):

Plug in '3y - 2' for 'x' in the first equation: x - 4y = 5 3y - 2 - 4y = 5 -y - 2 = 5 Add 2 to both sides, what's 5 + 2?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

Yes, so we have: -y = 7 Divide -1 to both sides, what's 7 / -1?

OpenStudy (anonymous):

wait so the answer is 7?

OpenStudy (anonymous):

No, we're not done yet.

OpenStudy (anonymous):

ohh ok '

OpenStudy (anonymous):

What's 7 / -1?

OpenStudy (anonymous):

um is it 6?

OpenStudy (anonymous):

No, it's -7. So we have: y = -7

OpenStudy (anonymous):

ohh ok thnk you sooooo much ...wait is that the answer?